# Learning-based Multi-View Stereo: A Survey

### v2

@@ 3.8.2 Offline MVS
$L=\sum_{\mathbf{p}}\left(\sum_{i=1}^{D}-G(i,\mathbf{p})\cdot\log[P(i,\mathbf{p})]\right),$ (11)
where $G(i,\mathbf{p})$ is the ground truth one-hot vector of depth at pixel $\mathbf{p}$, and $P(i,\mathbf{p})$ is the predicted depth probability.
@@
4 Unsupervised & Semi-supervised Methods with Depth Estimation
In this section, we introduce unsupervised learning-based MVS methods [125, 126, 42, 43, 41] and semi-supervised method [45]. Supervised MVS methods mentioned in Sec. 3 depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment. To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised learning-based methods that can provide competitive accuracy compared to the supervised ones without any ground truth depth. Existing unsupervised methods are built upon the assumption of photometric consistency (Sec. 4.1), and are categorized into end-to-end (Sec. 4.2) and multi-stage (Sec. 4.3). SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec. 4.4.
@@ 4.1 Photometric Consistency Assumption
In the realm of unsupervised depth map prediction, extant methods [127, 43, 128, 129] endeavor to establish photometric consistency between reference and source perspectives. This pivotal notion revolves around the augmentation of similarity between the reference image $\mathbf{I}_{0}$ and individual source images $\mathbf{I}_{i}$ after their warping to align with the reference view.
In relation to the depth estimation denoted as $\hat{d}$ for the initial image $\mathbf{I}_{0}$, the process involves the projection of reference pixels into the subsequent image $\mathbf{I}_{i}$ using the formulation presented in Eq. 4.
Subsequently, the distorted version of the source image denoted as $\hat{\mathbf{I}}_{0}^{i}$ is created by interpolating the RGB values of the source image at the displaced pixel positions through bilinear sampling. This interpolation is carried out at the locations where the pixels have been transformed due to the warping process. Additionally, alongside the warped image $\hat{\mathbf{I}}_{0}^{i}$, a binary mask $M_{i}$ is commonly generated. This mask is employed to exclude pixels that have been projected beyond the boundaries of the image and are, thus, considered invalid.
@@
$\begin{split}L_{\textrm{PC}}=&\sum_{i=1}^{N-1}\frac{1}{\left\|M_{i}\right\|_{1}}(\left\|\left(\hat{\mathbf{I}}_{0}^{i}-\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}\\ &+\left\|\left(\nabla\hat{\mathbf{I}}_{0}^{i}-\nabla\mathbf{I}_{0}\right)\odot M_{i}\right\|_{2}),\end{split}$ (12)
where $\nabla$ denotes the gradient at the pixel level, while $\odot$ symbolizes element-wise Hadamard multiplication. In most cases [127, 41, 43, 128, 42], the incorporation of structural similarity loss and depth smoothness loss into the computation is commonplace. This practice aims to enhance the training process’s stability and speed up the convergence.
The computation of the structural similarity loss occurs between a synthesized image and the initial reference image, aiming to uphold contextual congruity. More precisely, the assessment of contextual similarity often involves the Structural Similarity Index (SSIM) proposed [130], defined as:
@@
The incorporation of a smoothness loss term serves to promote the continuity of depth information within the context of image and depth disparity alignment. This continuity is evaluated based on the color intensity gradient present in the input reference image. The computation of the smoothness loss, $L_{\textrm{SM}}$, is defined as follows:
@@ 4.2 End-to-end Unsupervised Methods
Unsup_MVS [127] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps. However, photometric correspondences can be inaccurate due to non-Lambertian surfaces, camera exposure variations, and occlusions, leading to ambiguous supervision (Sec. 4.1). JDACS [41] introduces semantic consistency to address these challenges, using a pre-trained network to generate semantic maps and enforcing cross-view segmentation consistency. However, this approach can struggle to converge and lacks detailed information. RC-MVSNet [42] enhances supervision through neural rendering by combining Neural Radiance Fields and cost volumes. It uses a CasMVSNet backbone to generate depth maps supervised by photometric consistency, along with rendering-consistency and view synthesis losses to handle occlusions and varying lighting. ElasticMVS [131] addresses issues in photometric loss-based geometry by introducing a part-aware patch-match framework, using an elastic part representation to guide the depth map prediction. The network is optimized with contrastive and spatial concentration losses to promote pixel isolation. CL-MVSNet [132] improves proximity between positive pairs through contrastive consistency between a regular CasMVSNet branch and two contrastive branches. It also introduces $L0.5$ photometric consistency to improve the precision of the reconstruction.
@@ 4.3 Multi-stage Unsupervised Methods
Multi-stage methods require pre-training or data pre-processing and are based on pseudo-label generation.
Self-supervised CVP-MVSNet [128] generates pseudo ground truth depth using photometric consistency, refining depth maps with cross-view consistency checks and point cloud fusion [15]. The meshes are then reconstructed for further training iterations. U-MVS [43] pre-trains an optical-flow network, PWC-Net [133], and uses flow-depth consistency to generate pseudo labels, reducing supervision ambiguity in foregrounds and backgrounds with an uncertainty-aware self-training consistency. KD-MVS [44] employs knowledge distillation to increase performance. A teacher MVS model generates pseudo ground-truth labels through photometric and feature-metric consistency loss, with uncertainty encoding. These labels are then used to train the student models.
@@ 4.4 Semi-supervised Methods
5 Learning-based MVS without Depth Estimation
Though the learning-based methods that predict individual depth maps with plane-sweep are the main family of learning-based multi-view stereo, there are many methods of other families that achieve impressive 3D reconstruction quality in the recent years. In this section, We discuss four main families: voxel-based, NeRF-based, 3D Gaussian Splatting-based and large feed-forward methods.
@@ 5.1 Voxel-based Methods
These methods [46, 47, 134, 135] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF. Specifically, Atlas [46] and NeuralRecon [47] attempt to predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features. Atlas uses 3D CNN to regress the TSDF volume based on the feature volume accumulated from all images of the scene, which exhibits great completeness of reconstruction. The TSDF reconstruction is supervised using $L_{1}$ loss to the ground truth TSDF vaules. Since dense feature volume brings lots of computational overhead, NeuralRecon further improves the efficiency by incrementally reconstructing the scene in a fragment-wise and coarse-to-fine manner. The 3D features from different fragments are passed through the whole incremental reconstruction process with a RNN. TransformerFusion [134] fuses coarse and fine image features in a voxel grid with two transformers and then predicts an occupancy field to represent the scene geometry. VoRTX [135] uses a similar design to TransformerFusion and the scene geometry is obtained by passing three different-level features output from transformers through a 3D CNN.
@@ 5.2 NeRF-based Methods
$\mathcal{L}_{\text{color}}=\mathbb{E}[(||\mathbf{C}-\mathbf{C}_{gt}||^{2}].$ (17)
@@
Many subsequent endeavors [136, 137, 138, 139, 140, 141, 142, 143, 144] further improve NeRF in quality, fast training, memory efficiency and real-time rendering.
@@
Though the initial purpose of NeRF is to perform novel view synthesis, VolSDF [48] and NeuS [49] integrate NeRF with SDF for surface reconstruction. The SDF, denoted as $f$, is transformed into the density $\sigma$ for volume rendering. Specifically, for a point $\mathbf{p}(t)$, VolSDF computes the volume density $\sigma(\mathbf{p}(t))$ from the signed distance $f(\mathbf{p}(t))$ as:
$\sigma(\mathbf{p}(t))=\frac{1}{\beta}\Psi_{\beta}(-f(\mathbf{p}(t))),$ (18)
where $\Psi_{\beta}$ denotes the Cumulative Distribution Function (CDF) of a zero-mean Laplace distribution with learnable scale parameter $\beta>0$. NeuS computes the density in a different way as:
$\sigma(\mathbf{p}(t))=\max(\frac{-f^{\prime}(\mathbf{p}(t))\Phi^{\prime}_{s}(f(\mathbf{p}(t)))}{\Phi_{s}(f(\mathbf{p}(t)))},0),$ (19)
where $\Phi_{s}$ is the sigmoid function with learnable scale parameter $s$. After training, the mesh can be extracted from the SDF field with Marching Cubes [67].
Note that without ground truth geometry supervision, e.g., depth map or TSDF volume, these methods are trained in a self-supervised way as NeRF, Eq. 17. Since training with rendering loss only like NeRF has ambiguity in geometry [145], some following methods improve the reconstruction with explicit geometry supervision, e.g., monocular depth/normal priors [146, 147] and sparse SfM point cloud [50]. Motivated by Instant-NGP [138] that accelerates training with hash grids, other methods [148, 149, 150, 151] use hash grids to speed up training and improve surface details. In addition, to overcome the drawback that NeRF typically needs to be trained for each new scene, some recent methods [152, 153, 154] propose generalizable pipelines for implicit reconstruction even under sparse-view settings. Specifically, these methods project 3D points on the image planes and aggregate the corresponding image features as a feature volume like Atlas [46], which is used to estimate the surface location.
@@
Recall that the depth map-based MVS methods predict depth with photometric consistency across multiple views. However, this assumption fails for glossy surfaces with reflections and thus they cannot reconstruct them accurately. In contrast, some NeRF methods can handle reflections well. Recently, Ref-NeRF [155] reparameterizes the appearance with separate diffuse and reflective components by using the reflected view direction, which improves the rendering of specular surfaces. Therefore, recent methods [156, 157, 158, 159, 151] adopt this representation in reconstruction and can successfully reconstruct specular surfaces. Specifically, based on VolSDF/NeuS, these methods mainly replace the view direction with the reflected view direction following Ref-NeRF.
@@
where $x$ is the pixel location, $N$ is the number of sorted 2D Gaussians. During training, 3DGS minimizes the rendering loss like NeRF, as in Eq. 17. 3DGS can be initialized with SfM [161, 17] or MVS [33], which performs better than random initialization [160].
Motivated by the NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction. Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [51]. SuGaR [51] introduces a geometry regularization term encouraging the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better scene geometry. To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [162] on these points. NeuSG [163] jointly trains Neuralangelo [149], a NeRF-based method, and 3DGS. During training, the normals from the SDF field of Neuralangelo regularize the rotation of 3D Gaussians, while the SDF field is regularized to ensure that the SDF values at 3D Gaussians’ positions are close to zero. 2DGS [52] further improves surface accuracy by modeling the 3D primitives as surfels ( tangent plane clipped) and fuses rendered depth maps into mesh with TSDF fusion. PGSR [53] elicits unbiased depth maps as strong geometry prior from rendered distances and rendered normals, and enforces multi-view geometric consistency by regularizing such a prior with local plane assumption and edge constraint. RaDe-GS [164] also improves the quality of geometry by precisely modeling the depth and normal, instead of the error-prone approximate projection in the original 3DGS. However, retrieving good geometry from Gaussian-based representations still needs to be explored.
5.4 Large Feed-forward Methods
In 3D reconstruction and 3D generation, a recent prominent trend is to directly learn the 3D representation from large-scale 3D datasets, e.g.Objaverse [29]. These methods typically adopt large-scale transformers [165] to generate a 3D representation based on various inputs, including single image [54, 166, 167], textual description [168, 169], posed multi-view images [168, 170, 171, 172, 173], and un-posed multi-view images [174, 55]. LRM [54] uses a transformer-based feed-forward network to regress the 3D object from a single image. Specifically, it predicts features of a tri-plane representation [175], which is subsequently converted into a radiance field similar to NeRF. Large feed-forward models can also be combined with 3D Gaussian Splatting, where they directly generate 3D Gaussian points from posed multi-view inputs [168, 170, 171, 172, 173, 176]. Instead of predicting 3D Gaussian locations directly, which is a highly ill-posed task, the common practice is to use per-pixel aligned Plücker ray embeddings and predict the depth for each pixel [176] of input views. As such, this design requires the input views to cover a large range of the object. Recently, DUSt3R [55] predicts pixel-aligned point coordinates directly from uncalibrated image pairs. After extracting image token representations with Vision Transformer [177], two transformer decoders exchange information of the token representations via cross-attention and then regress the per-pixel 3D points in the coordinate frame of the first image. Many downstream tasks can be performed with the point clouds, e.g., point matching, localization, intrinsic and relative pose estimation. Directly extended from DUSt3R, MASt3R [56] focuses on the specific application of feature description and matching across multi-view images. It leverages an additional matching head on top of DUSt3R’s architecture to boost the performance of feature-based matching. With the dense correspondences established, MASt3R manages to achieve a comparable reconstruction quality as MVSNet [15]. Spann3R [178] manages an external spatial memory to keep track of all previous relevant 3D information yielded by DUSt3R and then queries this spatial memory to predict the 3D structure.

### v3

@@ 3.8.2 Offline MVS
$L=\sum_{\mathbf{p}}\left(\sum_{i=1}^{D}-G(i,\mathbf{p})\cdot\log[\mathbf{P}(\mathbf{p},i)]\right),$ (14)
where $G(\mathbf{p},i)$ is the ground truth one-hot vector of depth at pixel $\mathbf{p}$, and $\mathbf{P}(\mathbf{p},i)$ is the predicted depth probability.
@@
4 Unsupervised & Semi-supervised MVS with Depth Representation
In this section, we introduce unsupervised and semi-supervised MVS with depth representation. Supervised MVS methods mentioned in Sec. 3 depend extensively on the availability of accurate ground truth depth maps obtained through depth-sensing equipment. To make MVS practical in more general real-world scenarios, it is vital to consider alternative unsupervised methods that can provide competitive accuracy compared to the supervised ones without any ground truth. Existing unsupervised methods are built on the assumption of photometric consistency (Sec. 4.1), and are categorized into end-to-end (Sec. 4.2) and multi-stage (Sec. 4.3). To the best of our knowledge, SGT-MVSNet [45] is the only semi-supervised method so far, and we introduce it in Sec. 4.4.
@@ 4.1 Photometric Consistency Assumption
In the realm of unsupervised depth map prediction, extant methods [135, 43, 136, 137] endeavor to establish photometric consistency between reference and source views. This pivotal notion revolves around the augmentation of similarity between the reference image $\mathbf{I}_{0}$ and source image $\mathbf{I}_{i}$ after warping to align with the reference view. Specifically, for each pixel $\mathbf{p}_{0}$ in $\mathbf{I}_{0}$, we use the estimated depth value $\hat{d}(\mathbf{p}_{0})$ to project $\mathbf{p}_{0}$ into $\mathbf{I}_{i}$ with intrinsic and extrinsic matrices, following Eq. 5. To get the warped source image $\hat{\mathbf{I}}_{0}^{i}$, we interpolate the RGB values from the source image $\mathbf{I}_{i}$ at the coordinates $\mathbf{p}_{i}$ with bilinear interpolation. Additionally, alongside the warped image $\hat{\mathbf{I}}_{0}^{i}$, a binary mask $M_{i}$ is commonly generated. This mask is employed to exclude pixels that have been projected beyond the boundaries of the image and are, thus, considered invalid.
@@
$\begin{split}L_{\textrm{PC}}=&\sum_{i=1}^{N-1}\frac{1}{\left\|M_{i}\right\|_{1}}(\left\|(\hat{\mathbf{I}}_{0}^{i}-\mathbf{I}_{0})\odot M_{i}\right\|_{2}\\ &+\left\|(\nabla\hat{\mathbf{I}}_{0}^{i}-\nabla\mathbf{I}_{0})\odot M_{i}\right\|_{2}),\end{split}$ (15)
where $\nabla$ denotes the gradient at the pixel level, while $\odot$ symbolizes element-wise Hadamard multiplication.
In most cases [135, 41, 43, 136, 42], the incorporation of structural similarity loss, $L_{\textrm{SSIM}}$, and depth smoothness loss, $L_{\textrm{SM}}$, into the computation is common, which improves the stability of training process and speeds up the convergence. Based on the Structural Similarity Index (SSIM) [138], structural similarity loss is computed between a synthesized image and the reference image as:
@@
The incorporation of a smoothness loss term serves to promote the continuity of depth information within the context of image and depth disparity alignment. This continuity is evaluated based on the color intensity gradient present in the input reference image. The smoothness loss, $L_{\textrm{SM}}$, is defined as:
@@ 4.2 End-to-end Unsupervised Methods
Khot et al. [135] adopts view synthesis supervision and dynamically selects the $X$ best (lowest loss) values out of $Y$ loss maps. However, photometric correspondences can be inaccurate due to non-Lambertian surfaces, camera exposure variations, and occlusions, leading to ambiguous supervision (Sec. 4.1). JDACS [41] introduces semantic consistency to address these challenges, using a pre-trained network to generate semantic maps and enforcing cross-view segmentation consistency. However, this approach can struggle to converge and lacks detailed information. RC-MVSNet [42] enhances supervision through neural rendering by combining NeRF [28] and cost volumes. Based on a CasMVSNet [32] backone that is supervised by photometric consistency, it uses rendering consistency and view synthesis loss to handle occlusions and varying lighting. ElasticMVS [139] addresses issues in photometric loss-based geometry by introducing a part-aware patch-match framework, using an elastic part representation to guide the depth map prediction. The network is optimized with contrastive and spatial concentration losses to promote pixel isolation. CL-MVSNet [140] improves proximity between positive pairs through contrastive consistency between a regular CasMVSNet branch and two contrastive branches. It also introduces $L_{0.5}$ photometric consistency loss, which is more robust to outliers because it penalizes large errors less severely than traditional $L_{1}$ or $L_{2}$ norms. By growing sub-linearly with the error, it prevents the model’s training from being dominated by outliers with large errors. This allows the network to focus on well-matched regions.
@@ 4.3 Multi-stage Unsupervised Methods
Multi-stage methods require pre-training or data pre-processing and are typically based on pseudo-label generation. Self-supervised CVP-MVSNet [136] generates pseudo ground truth depth using photometric consistency, geometric consistency and point cloud fusion [9]. U-MVS [43] pre-trains an optical-flow network, PWC-Net [141], and uses flow-depth consistency to generate pseudo labels, reducing supervision ambiguity in foregrounds and backgrounds with an uncertainty-aware self-training consistency. KD-MVS [44] employs knowledge distillation to increase performance. A teacher MVS model generates pseudo ground-truth labels through photometric and feature-metric consistency loss, with uncertainty encoding. These labels are then used to train the student models.
@@ 4.4 Semi-supervised Methods
5 Learning-based MVS with Other Representations
n this section, we discuss learning-based MVS with the following representations: voxel grids, NeRF, 3D Gaussian Splatting, and direct point-based representations. These approaches have demonstrated impressive 3D reconstruction capability in recent years.
@@ 5.1 Voxel-based Methods
These models [46, 47, 48, 142] estimate the scene geometry with volumetric representation by leveraging implicit function, e.g., SDF. Specifically, these methods predict the TSDF volume from the 3D feature volume constructed by lifting 2D image features and extract the mesh with marching cube [73].
5.1.1 Direct 3D CNN
As a seminal work, Atlas [46] employs a 3D CNN with four scale resolution pyramid to regress the TSDF volume from the feature volume accumulated from all images, as shown in Fig. 4. The TSDF reconstructions are supervised using $L_{1}$ loss to the ground truth TSDF vaules. However, constructing a dense 3D feature volume is a huge computational overhead for large scenes.
5.1.2 Coarse-to-fine
It is a common practice to adopt a coarse-to-fine paradigm [47, 48, 142] for the sake of efficiency. One of its core steps is voxel sparsification, which eliminates empty voxels at the coarse level so that memory consumption is reduced in the fine level. NeuralRecon [47] incrementally reconstructs the scene in a coarse-to-fine and fragment-wise manner. First, in order to reduce memory, NeuralRecon predicts TSDF with a three-level coarse-to-fine approach that gradually increases the density of sparse voxels. Second, it reconstructs the TSDF volume of a local fragment and then fuses it with the global volume by GRU [143] Fusion.
Based on the coarse-to-fine framework of NeuralRecon, TransformerFusion [48] introduces transformer [144] that enables the network to learn to attend to the most relevant image frames for each 3D location in the scene. Similarly, VoRTX [142] uses a transformer to fuse image features at each voxel to produce the multi-view feature volume.
Although state-of-the-art voxel-based methods use coarse-to-fine framework to explicitly improve the efficiency, they are still limited to indoor scenes that are strictly bounded, e.g., ScanNet [27], since the memory storage explicitly increases when the scale of scene increases, e.g., unbounded outdoor scenes.
Figure 4: Pipeline of Atlas [46]. 2D image features are back-projected into 3D volumes, which are aggregated and passed through a 3D CNN to directly regress a TSDF volume.
@@ 5.2 NeRF-based Methods
[…]
@@
The pipeline of NeRF is shown in Fig. 5. Many subsequent endeavors [145, 146, 147, 148, 149, 150, 151, 152, 153] further improve NeRF in quality, fast training, memory efficiency and real-time rendering.
Figure 5: Pipeline of NeRF [28]. Given a 3D position and 2D viewing direction (a), an MLP produces the color and volume density (b). Then volume rendering is used to composite these values into an image (c). The optimization is minimizing the rendering loss (d).
Figure 6: Pipeline of 3DGS [29]. 3DGS initializes 3D Gaussian primitives from scene geometry priors (e.g., SfM points) and computes pixel colors by compositing their projected attributes via 3D-to-2D splatting.
Figure 7: Pipeline of DUSt3R [61]. Two views of a scene $(I^{1},I^{2})$ are first encoded in a Siamese manner with a shared ViT encoder. The resulting token representations $F^{1}$ and $F^{2}$ are then passed to two transformer decoders that constantly exchange information via cross-attention. Finally, two regression heads output the two corresponding pointmaps and associated confidence maps. Importantly, the two pointmaps are expressed in the same coordinate frame of the first image $I^{1}$.
@@
Though the initial purpose of NeRF is to perform novel view synthesis, VolSDF [154] and NeuS [49] integrate NeRF with SDF for surface reconstruction. The SDF, denoted as $f$, is transformed into the density $\sigma$ for volume rendering. After training, the mesh can be extracted from the SDF field with Marching Cubes [73]. Since the main focus of this survey is 3D reconstruction, we discuss these NeRF-based methods with SDF representation in the following.
5.2.1 Optimization-based NeRF
Without ground truth geometry supervision, e.g., depth map or TSDF volume, training is mainly done in a self-supervised manner with rendering loss, Eq. 20. To enforce SDF function $f$ approximate a valid SDF, existing methods [154, 49] use an eikonal loss [155] for regularization:
[…]
where $\mathbf{x}$ denotes the 3D point. Motivated by Instant-NGP [147] that accelerates training with hash grids, many methods [156, 50, 157, 51] use hash grids to speed up training and improve surface details.
However, these methods still mainly use pixel-wise rendering loss for training, which may result in geometric ambiguity, especially in textureless areas [158]. To alleviate this issue, some methods [159, 160] use patch-wise loss to reduce ambiguity. Specifically, based on rendered depth and normal, they warp reference image patches to neighboring views and measure photometric consistency, which is similar to unsupervised depth map-based MVS. In addition, other methods introduce explicit geometry supervision, e.g., monocular depth/normal priors [161, 162] and sparse SfM point cloud [160]. For example, MonoSDF [161] uses a pre-trained Omnidata model [163] to predict monocular depth and normal, which are used to enforce consistency with rendered depth and normal. Recently, Ref-NeRF [164], a variant of original NeRF, reparameterizes the appearance with separate diffuse and reflective components by using the reflected view direction, which improves the rendering of specular surfaces. Therefore, some follow-up methods [165, 166, 167, 168, 51] adopt this representation in reconstruction and can successfully reconstruct specular surfaces.
@@ 5.2.1 Optimization-based NeRF
Recall that the depth map-based MVS methods predict depth with photometric consistency across multiple views. However, this assumption fails for glossy surfaces, and thus these methods cannot reconstruct them accurately. Therefore, we believe that NeRF-based MVS is more suitable than depth map-based MVS for reconstructing specular surfaces.
5.2.2 Generalizable NeRF
NeRF usually requires a lengthy optimization for each scene and cannot generalize to unseen scenes in a zero-shot manner. To address this issue, recent methods [169, 52, 53, 170] propose generalizable NeRF pipelines for both rendering and reconstruction. Interestingly, we observe that the majority of these methods are inspired by depth map-based and voxel-based MVS. For example, MVSNeRF [169] and C2F2NeUS [171] construct cost volumes with plane sweep to encode global features, following depth map-based offline MVS methods [9, 32]. Similar to voxel-based MVS [46], some methods [52, 53, 170] use image features to construct feature volumes for geometry estimation. More recently, large reconstruction models (LRMs) [54] leverage transformer backbones to directly predict NeRF representations from single or sparse images in a feed-forward manner. These models support inputs ranging from single image [54, 172, 173] to posed [174, 175] and un-posed multi-view imagery [176], significantly improving practicality beyond per-scene optimization.
@@ 5.2.2 Generalizable NeRF
where $x$ is the pixel location, $N$ is the number of sorted 2D Gaussians.
@@
During training, 3DGS minimizes the rendering loss like NeRF, as in Eq. 20. 3DGS can be initialized with SfM [178, 17] or MVS [30], which performs better than random initialization [177]. The pipeline of 3DGS is shown in Fig. 6.
5.3.1 Optimization-based 3DGS
Motivated by NeRF-based MVS methods, researchers try to adapt 3DGS for reconstruction task. Though 3DGS achieves high-quality novel-view synthesis, it is challenging to recover high quality geometry since no explicit geometry constraint is used and 3D Gaussians do not correspond well to the actual surface because of the 3D covariance [55].
In order to make 3D Gaussians correspond well to the actual surface, SuGaR [55] introduces a geometry regularization term that encourages the 3D Gaussians to be well-aligned over the scene surfaces so that the Gaussians can contribute to better geometry. To reconstruct the mesh, SuGaR samples 3D points on a level set of the density computed from 3D Gaussians and then runs Poisson Reconstruction [179] on these points. Instead of performing regularization on 3D Gaussian, 2DGS [56] introduces a 2D Gaussian representation by modeling the 3D primitives as surfels.
Similar to NeRF, relying solely on image rendering loss can easily fall into local minima, leading to Gaussian shapes inconsistent with the actual surface. Introducing explicit geometry constraints can alleviate this problem. For example, PGSR [57] enforces the local consistency of depth and normal in a local image patch, and multi-view geometric and photometric consistency as COLMAP [17].
5.3.2 Generalizable 3DGS
Following the paradigm of feed-forward NeRF, PixelSplat [180] predicts 3DGS representations from paired images by leveraging epipolar constraints to resolve scale ambiguity. MVSplat [58] further incorporates depth cues for 3DGS prediction by constructing a cost volume from multiple input images. DepthSplat [59] extends this idea by achieving more robust depth estimation through the combination of cost volume and monocular depth features [181]. Instead of focusing on improving geometry and rendering quality, recent works such as SplatterImage [182], Flash3D [183], GS-LRM [184],LGM [60], and Bolt3D [185] explore the generalization capability of single-image-to-3DGS generation without per-scene optimization. These methods leverage multi-view diffusion models or transformers pretrained on large-scale multi-view datasets to directly predict per-pixel 3D Gaussian attributes from images.
5.4 Large Feed-forward Point-based Methods
In 3D reconstruction, a prominent trend is to directly learn 3D point geometry from large-scale multi-view data with transformer backbones [144], shifting the paradigm from per-scene optimization toward one-shot inference of geometry or renderable content. Unlike traditional MVS that relies on hand-crafted cost volumes, these methods shift the paradigm toward one-shot inference of dense point maps.
Recent feed-forward point-based methods compress the classical MVS pipeline into a single forward pass: (1) feature extraction; (2) cross-view aggregation via attention in lieu of an explicit plane-sweep cost volume; and (3) prediction heads that jointly regress cameras and dense geometry (depth/point maps). We illustrate the pipeline of DUSt3R [61], a pioneering feed-forward method, in Fig. 7. Below we describe how modern feed-forward methods instantiate each stage in a single forward pass.
ViT feature extraction. A pre-trained transformer encoder extracts view-wise, geometry-aware features/tokens from few to many images, supporting downstream camera, depth, and pointmap prediction [62, 61, 186, 63]. In practice, DINO models are often chosen as the ViT backbone because its self-supervised features transfer robustly across tasks and domains and exhibit emergent semantics useful for dense correspondences [187].
Cross-view aggregation via attention. Instead of building explicit plane-sweep cost volumes, global or alternating attention patterns interleave per-frame and multi-frame aggregation, allowing long-range correspondence reasoning while remaining feed-forward [62, 61]. The attention mechanism functions as a learnable and unconstrained feature matching in multi-view feed-forward 3D reconstruction. With no pose prior and unknown epipolar lines, the theoretical area to be searched is actually the whole image and the Transformer modules learn implicit correspondences in a data-driven manner, as is revealed in [188]. While the more classical cost volumes are built with known poses and epipolar lines, where the area to be searched become a 1D segment.
Prediction heads for geometry and pose. Rather than deriving geometry purely from pairwise matches, modern heads directly regress intrinsics/extrinsics together with dense depth or point maps. Early work showed that regressing point maps is effective even from two views and unifies monocular/binocular settings [61]; recent multi-view extensions output cameras, depth, point maps, and tracks in one pass [62]. Emerging designs relax reference-view assumptions via permutation-equivariant architectures and scale to long sequences with spatial memory or causal/streaming transformers [189, 190, 191].
For future work, our analysis points to two avenues for advancing feed-forward 3D reconstruction. First, to address the fundamental out-of-distribution problem, future research may focus on developing memory-efficient training paradigms that allow large-scale, many-view scenes to be incorporated directly during training. By aligning the training distribution more closely with real-world test scenarios, such approaches could substantially improve generalization to complex environments. Second, to overcome the accuracy limitations of single-pass prediction, a promising direction is to create hybrid architectures that merge feed-forward speed with iterative refinement.